OpenStudy (anonymous):

The sum of a man's age and his grandmother's age is 102. Eighteen years ago his grandmother was 10 times as old as he was then. Find their present ages.

OpenStudy (anonymous):

Let man's age be: x And grandmother's age be: y x + y = 102 y = 102 - x (y - 18) = 10(x - 18) 102 - x -18 = 10x - 180 11x = 264 x = 24 So y = 102 - 24 y = 78..